Three forces have the following components, all in Newton’s: (Fx = -5, Fy = -7), (Fx = 0, Fy = 30),
(Fx 20, Fy = -16). Find their resultant.
1
Expert's answer
2021-09-14T09:35:37-0400

The resultant can be found by Pythagorean theorem:


"F_R=\sqrt{F_{xR}^2+F_{yR}^2}=\\
=\sqrt{(-5+0+20)^2+(-7+30-16)^2}=16.6\text{ N}."